OpenStudy (anonymous):

Solve b2 + 8b – 30 = 3 by completing the square

OpenStudy (aripotta):

i'm not quite sure how to do this :/

OpenStudy (aripotta):

is there multiple choice?

jimthompson5910 (jim_thompson5910):

\[\Large b^2 + 8b - 30 = 3\] \[\Large b^2 + 8b = 3+30\] \[\Large b^2 + 8b = 33\] \[\Large b^2 + 8b+16 = 33+16\] \[\Large (b+4)^2 = 49\] See where to go from here?

jimthompson5910 (jim_thompson5910):

In the second to the last step I took half of 8 to get 4. Then squared 4 to get 16. From there, I added 16 to both sides.

jimthompson5910 (jim_thompson5910):

Keep going to get \[\Large (b+4)^2 = 49\] \[\Large b+4 = \pm\sqrt{49}\] \[\Large b+4 = \sqrt{49} \ \text{or} \ b+4 = -\sqrt{49}\] \[\Large b+4 = 7 \ \text{or} \ b+4 = -7\] \[\Large b = -4+7 \ \text{or} \ b = -4-7\] \[\Large b = 3 \ \text{or} \ b = -11\]